# A coupled finite element-virtual element method for thermomechanical analysis of electronic packaging structures

### v2

@@ 3.1.1 Thermal analysis
For the finite element discretization of the thermal problem, we employ the weak formulation established in Sec. $2.$ The domain is discretized into finite elements such that $\Omega\equiv\text{int}\left(\bigcup_{i}E_{i}\right)$, where the operator ‘int’ denotes the interior of a set, $\Omega$ represents the union of the interior regions of all elements $E_{i}$, excluding the boundaries. And element Ei is chosen as a four-node quadrilateral isoparametric element.
@@ 3.1.2 Thermoelastic analysis
For the finite element discretization of the mechanical problem, we employ the weak formulation established in Sec. $2.$ Within each element, the displacement field is interpolated using the same shape functions as the thermal analysis
@@ 4.3.2 Thermal analysis phase
The thermal analysis phase implements the heat conduction formulation established in Sec. $2.$ For the FE domain, standard shape functions are constructed and element thermal stiffness matrices $\mathbf{K}_{\mathrm{FE}}^{th,E}$ are computed using the procedures described in Sec. 3.1.1. Simultaneously, the VE domain employs projection operators to calculate element stiffness matrices $\mathbf{K}_{\mathrm{VE}}^{th,E}$ following the VEM formulation.
@@ 4.3.3 Thermoelastic analysis phase
The mechanical analysis phase utilizes the computed temperature field as thermal loading input, implementing the thermomechanical coupling described in Sec. $2.$ Following procedures analogous to the thermal analysis, shape functions are constructed and element mechanical stiffness matrices are computed for both domains: $\mathbf{K}_{\mathrm{FE}}^{u,E}$ for the finite element region and $\mathbf{K}_{\mathrm{VE}}^{u,E}$ for the virtual element region.
@@ 5.1 Square plate with circular hole
where NN represents the total number of nodes in the calculation, $\bm{\sigma}_{\text{num}}$ denotes the numerical solution of nodal stress, and $\bm{\sigma}_{\text{ref}}$ represents the reference solution of nodal stress obtained from mesh refinement (nDof =334,254) in Abaqus.
@@ 5.2 Thermoelastic analysis of thick-walled cylindrical structure
where NN is the number of evaluation points, and $X_{e}$ and $X$ denote the analytical and numerical solutions, respectively.
@@ 5.3 Simplified sintered silver model
The third numerical example investigates a sintered silver interconnect structure commonly employed in high-power electronic packaging applications. As illustrated in Fig. 10, the model adopts a typical ‘sandwich’ configuration consisting of three distinct layers: a SiC chip (top), sintered silver interconnect layer (middle), and copper substrate (bottom). This multilayer structure is representative of modern power electronic modules where effective thermal management and mechanical reliability are critical design considerations [yu2021thermal]. The geometric dimensions are specified as follows: the SiC chip has width $b=1.8$ mm and thickness $h_{1}=0.5$ $m$; the sintered silver layer maintains the same width as the SiC chip with thickness $h_{2}=0.3$ $m$; and the copper substrate has width $a=3$ mm and thickness $h_{3}=0.8$ $m$. The material properties for each component are detailed in Table 3, which reveal significant differences in thermal and mechanical characteristics that contribute to complex thermomechanical stress development during thermal loading.

### v3

@@ 3.1.1 Thermal analysis
For the finite element discretization of the thermal problem, we employ the weak formulation established in Sec. 2.2. The domain is discretized into finite elements such that $\Omega\equiv\text{int}\left(\bigcup_{i}E_{i}\right)$, where the operator ‘int’ denotes the interior of a set, $\Omega$ represents the union of the interior regions of all elements $E_{i}$, excluding the boundaries. And element Ei is chosen as a four-node quadrilateral isoparametric element.
@@ 3.1.2 Thermoelastic analysis
For the finite element discretization of the mechanical problem, we employ the weak formulation established in Sec. 2.2. Within each element, the displacement field is interpolated using the same shape functions as the thermal analysis
@@ 4.3.2 Thermal analysis phase
The thermal analysis phase implements the heat conduction formulation established in Sec. 2.2. For the FE domain, standard shape functions are constructed and element thermal stiffness matrices $\mathbf{K}_{\mathrm{FE}}^{th,E}$ are computed using the procedures described in Sec. 3.1.1. Simultaneously, the VE domain employs projection operators to calculate element stiffness matrices $\mathbf{K}_{\mathrm{VE}}^{th,E}$ following the VEM formulation.
@@ 4.3.3 Thermoelastic analysis phase
The mechanical analysis phase utilizes the computed temperature field as thermal loading input, implementing the thermomechanical coupling described in Sec. 2.2. Following procedures analogous to the thermal analysis, shape functions are constructed and element mechanical stiffness matrices are computed for both domains: $\mathbf{K}_{\mathrm{FE}}^{u,E}$ for the finite element region and $\mathbf{K}_{\mathrm{VE}}^{u,E}$ for the virtual element region.
@@ 5.1 Square plate with circular hole
where $N$ represents the total number of nodes in the calculation, $\bm{\sigma}_{\text{num}}$ denotes the numerical solution of nodal stress, and $\bm{\sigma}_{\text{ref}}$ represents the reference solution of nodal stress obtained from mesh refinement (nDof =334,254) in Abaqus.
@@ 5.2 Thermoelastic analysis of thick-walled cylindrical structure
where $N$ is the number of evaluation points, and $X_{e}$ and $X$ denote the analytical and numerical solutions, respectively.
@@ 5.3 Simplified sintered silver model
The third numerical example investigates a sintered silver interconnect structure commonly employed in high-power electronic packaging applications. As illustrated in Fig. 10, the model adopts a typical ‘sandwich’ configuration consisting of three distinct layers: a SiC chip (top), sintered silver interconnect layer (middle), and copper substrate (bottom). This multilayer structure is representative of modern power electronic modules where effective thermal management and mechanical reliability are critical design considerations [yu2021thermal]. The geometric dimensions are specified as follows: the SiC chip has width $b=1.8$ mm and thickness $h_{1}=0.5$ mm; the sintered silver layer maintains the same width as the SiC chip with thickness $h_{2}=0.3$ mm; and the copper substrate has width $a=3$ mm and thickness $h_{3}=0.8$ mm. The material properties for each component are detailed in Table 3, which reveal significant differences in thermal and mechanical characteristics that contribute to complex thermomechanical stress development during thermal loading.